A flat coil enclosing an area of 0.18 m2 is rotating at 65 rev/s, with its axis of rotation perpendicular to a 0.22 T magnetic field.

1. If there are 1075 turns on the coil, what is the maximum voltage induced in the coil?

2. When the maximum induced voltage occurs, what is the orientation of the coil with respect to the magnetic field?

a. The plane of the coil is perpendicular to the magnetic field.

b. The plane of the coil is at a 43° angle to the magnetic field.

c. The plane of the coil is parallel to the magnetic field.

d. The plane of the coil is at a 65° angle to the magnetic field.
Answer #1

Here we apply Faraday's law of electromagnetic induction and definition of magnetic flux.

ØB Area of flat coil, A=0.18m2 langular speed, w = 6S rey/sec - =65(20) radeel = 408.4 rad/aec | Magnetic field, B= 0.225 No| As do = w NBA Go Sin O Eman will be for sino = 1 B Emar = N BAW=1075)(0:22 (0.18 4682 » Eman =(1075) (22) (6.18) (408-41) wy